Cone
Height= 32 inches
Diameter= 11.5 inches
Can you create a problem in a form of algebraic expression in finding its surface area.? And please show me the solution. Thank you.

Answer by addingup(3677)

A = πr(r+√h^2+r^2)
In your problem: Diameter 11.5/2= 5.75 radius
A= π5.75(5.75√32^2+5.75^2)
A= 18.06(5.75√1057.06)
A= 18.06(5.75*32.51)
A= 18.06*186.93
A= 3,376